You decide that you need $88,000 in 10 years in order to make a down payment...

  1. You decide that you need $88,000 in 10 years in order to make a down payment on a house. You plan to make annual deposits to achieve your goal. If interest rate is 8%, how much should be deposited each time? Hint use the Financial Function “PMT” to solve for the payment and PV will be zero.

  1. You owe $46,000 to your parents for funding some of your college. You promise to make 8 annual payments of $8,000 to settle your debt. Approximately what interest rate are your parents charging (estimated up to 2 decimal places), if you make the 8 annual payments beginning one year from now? Hint use the Financial Function “RATE”. Note if you need help use the “Help on this function” feature. This help will include an example for you to follow.
  1. You are offered an investment that will pay $36,300 per year for 9 years. If you feel that the appropriate discount rate is 7%, what is the investment worth to you today?

  1. Your grandparents offered you some money. You are offered the following options. Assuming an annual interest rate of 5.0%, which option should you choose? Show your work for each alternative

  1. Receive $20,000 immediately
  2. Receive $3,200 at the end of each six months for four years. You will receive the first check in six months
  3. Receive $2,900 at the end of each year for four years, and then $20,000 at the end of the fifth year.
Answer #1

1) For PMT the calculation will be in Excel =PMT(8%/1,8*1,you want $88,000) =$8,273 you need to pay annually

2) APR Formula is APR=((Fees+Interest/Principal/No.of Days)*365)*100 so the calculation will be

((0+$18,000/$46,000/2920)*365)*100=4.89%Interest Rate

3) $36,300*9=$326,700 Total Payment ROI 7% for 9 years will be=$160,083

you need to pay $326,700-$160,083=$166,617

4) Scenario a) $20,000

Scenario b) $3,200*8=$25,600*5/100*4= $5,120+$25,600=$30,720

Scenario c) 4th year $2,900+$2,900*5%*4=$12,180

end of 5th year will get $20,00 so the total calculation after 5th year will be=$20,000+ ($20,000*5/100)=$21,000+$12,180=$33,180

So the C option will be more monetary benefited.